solve the system of equations using substitution 5x+y=32
y=3x

Answer by Alan3354(69443)

solve the system of equations using substitution 5x+y=32
y=3x
--------------
Sub is a good choice, since y is already solved for in the 2nd eqn.
Sub for y in the 1st eqn
------------
5x + y = 32
5x + 3x = 32
8x = 32
x = 4
------
y = 3x
y = 12